A random variable X has Poisson distribution with mean equal to 0.4. What is the probability that the random variable is greater than zero?.
1
Expert's answer
2020-06-04T21:17:35-0400

"P(X>0)=1-P(X=0)\\
=1-\frac{e^{-0.4}(0.4)^{0}}{0!}=1-0.67=0.33"